# Nova: Generative Language Models for Assembly Code with Hierarchical Attention and Contrastive Learning

### v5

@@ 4.2 Binary Code Similarity Detection
Tables 9,  9,  9 and  9 show the Recall@1 of Nova and existing BCSD techniques with pool size $K$ of 50, 100, 200 and 500 on the six benchmarks. Underline indicates the best in each benchmark, while wave denotes the tied best (we only mark Nova-1B for clearer illustration).
Overall, Tables 9,  9,  9 and 9 show that on average, Nova-1B and Nova-6B achieve the highest Recall@1 (in bold) under all four settings of $K$. Nova-6B further outperforms Nova-1B and achieves the highest averaged Recall@1 under all four settings, ranking the ground-truth of 5%, 2%, 4%, and 3% more queries the most similar correspondingly compared to CodeArt. Nova-1B consistently outperforms existing techniques with higher Recall@1 when $K$ is 50, 100, and 200, meaning it correctly ranks ground-truth of 3%, 1%, and 2% more queries as the most similar. Under the setting of $K=500$, Nova-1B ties with CodeArt with the same highest Recall@1. When looking into each individual benchmark, Nova-1B always wins on the most benchmarks under different settings of pool size $K$. For instance, Nova-1B wins on four benchmarks while DiEmph only wins on SQLite when $K=50$. We also conduct an ablation study on BCSD in Appendix A.5.
@@ 4.3 Analytic Experiments: How are Nova’s embeddings better?
Figure 5: t-SNE analysis of embeddings cauclated by Nova-CL-HA, Nova-HA, and Nova.
5 Related Work
5.1 Binary Models
Machine learning models are widely used in binary program analysis tasks. However, these models are typically designed for specific tasks such as binary code similarity detection (Pei et al., 2020; Xu et al., 2023a; Wang et al., 2022), variable name prediction (Chen et al., 2022c; Xu et al., 2023b; Zhang et al., 2021), binary code type inference (Pei et al., 2021), and are built from scratch.
5.2 Large Source-Code Models
LLMs demonstrate promising results on many code-related tasks, such as code generation (Chen et al., 2022a; Liu et al., 2023; Chen et al., 2023; Le et al., 2022; Yue et al., 2021; Chen et al., 2021; Nijkamp et al., 2022; Fried et al., 2023; Rozière et al., 2023; Guo et al., 2024; Lozhkov et al., 2024; Hui et al., 2024), bug fixing (Jiang et al., 2023; Xia et al., 2023) and vulnerability fixing (Wu et al., 2023; Steenhoek et al., 2023; He & Vechev, 2023). The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets (Touvron et al., 2023; OpenAI, 2023). Nova is designed and trained for assembly, which has unique challenges such as low information density and diverse optimization.
5.3 Attention Mechanism
Standard self-attention is widely used in transformer architecture (Vaswani et al., 2023) to capture soft dependencies between tokens in the input. Many special attention mechanisms have been designed for better learning in various scenarios (Yang et al., 2016; Huang et al., 2024). LongCoder (Guo et al., 2023) combines window attention and global attention (attention sink (Xiao et al., 2024)) to handle long input of source code. We have shown that LongCoder’s window attention is less effective than Nova’s on assembly code.
CAST (Shi et al., 2021) is a new neural architecture that splits the abstract syntax tree (AST) of source code into subtrees, encodes the subtrees, and aggregates to the final encoding. PA-former (Chai & Li, 2024) is a new neural architecture that constructs source code as pyramid input based on their AST structure and contains a pyramid attention mechanism to calculate the features in a hierarchical way. HierarchyNet (Nguyen et al., 2023) is a neural architecture that considers source code AST, data flow, and control flow graphs. Similarly, it cannot be applied to assembly code. Different from CAST, PA-former, and HierarchiyNet, Nova’s attention design is for assembly code, is more lightweight and can be plugged into any pre-trained generative LLM.
6 Limitations
One limitation is that Nova is X86-specific, as we only collect X86 assembly corpus for pre-training. This design choice is mainly affected by two facts: (1) X86 assembly is used and explored in a wide range of binary tasks (Wang et al., 2022; Su et al., 2024; Xu et al., 2023a; Chen et al., 2022c) compared to other assembly languages, and (2) computation limitations. However, the proposed techniques are independent of X86 assembly. Low information density and compiler optimization are the common challenges of most assembly languages such as X86, ARM, and MIPS. The proposed techniques can be applied to the future development of multi-lingual assembly LLMs.
Another potential limitation is the scale of models. We develop Nova-1B and Nova-6B. These two sized LLMs show impressive ability in assembly code decompilation and encoding. There might be potential benefit of developing larger Nova models. However, due to the computing resources limitation, we are unable to explore that in this work.
@@ 7 Conclusion
This work develops Nova, a generative foundation LLM for assembly code, which incorporates two key novelties (hierarchical attention and contrastive learning objectives) to address the unique challenges of assembly code. Evaluation on downstream tasks shows the effectiveness of Nova, which outperforms existing techniques on binary code decompilation by up to 146.54% and outperforms the latest binary code similarity detection techniques by up to 6.17%. We expect our hierarchical attention and contrastive learning techniques to benefit source code and natural language foundation models, which remains as future work.
@@ ACKNOWLEDGMENTS
This research was supported in part by NSF 1901242 and 2006688 and a CFI fund.
This paper was prepared for informational purposes in part by the CDAO group of JPMorgan Chase & Co and its affiliates (”J.P. Morgan”) and is not a product of the Research Department of J.P. Morgan. J.P. Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
@@ A.1 Data Collection
A.2 Training Details
This section provides additional details of training. We pre-train Nova starting from DeepSeek-Coder, using the language modeling objective ($L_{lm}$) for one epoch on the AnghaBench and The-Stack corpora. This step uses a batch size of 128, with the input truncated by a 1,024 tokens limit. The model weights are updated using the AdamW optimizer. The learning rate is $5e^{-5}$, using 1000 steps of warm-up and a cosine decay to adjust the learning rate.
@@ A.2 Training Details
Infrastructure The training are conducted on eight NVIDIA RTX A100 GPUs, each with 40GB memory. Our implementation is based on Huggingface’s implementation of DeepSeek-Coder  [^1] https://huggingface.co/deepseek-ai/deepseek-coder-1.3b-base, PyTorch  [^2] https://pytorch.org/get-started/locally/, and DeepSpeed  [^3] https://github.com/microsoft/DeepSpeed.
@@ A.3 Applying Hierarchical Attention on Half Attention Heads
Figure 7: Nova-1B correctly decompiles HumanEval-Decompile task 0.
@@
Figure 8: Nova-1B correctly decompiles HumanEval-Decompile task 20.
Benchmarks Nova-CL-HA Nova-HA Nova-1B
@@
Table 11: Ablation study with $K=50$.
Benchmarks Nova-CL-HA Nova-HA Nova-1B
@@
Table 12: Ablation study with $K=100$.
Benchmarks Nova-CL-HA Nova-HA Nova-1B
@@
Table 13: Ablation study with $K=200$.
Benchmarks Nova-CL-HA Nova-HA Nova-1B
@@
Table 14: Ablation study with $K=500$.
A.4 Binary Code Decompilation Case Studies
Figure 7 shows an example from HumanEval-Decompile (task_id 0). Given the O1-optimized assembly code, GPT-4 fails to figure out the number of function arguments correctly, missing one important argument “float e”, and thus produces wrong functionality in the decompiled code. LLM4Decompile-1B makes similar mistakes and also misses the inner nested for loop. Nova-1B correctly decompiles the assembly into source code, where the ground truth is checking if any two elements in the given list *a (with size n) are close to each other than a given threshold e.
Figure 8 shows another more complex example, HumanEval-Decompile task_id 20. Nova-1B correctly decompiles the source code, successfully figuring that the function is trying to find the two elements that are closest to each other in the given array *a, with the result stored in minmax.
@@ A.4 Binary Code Decompilation Case Studies
A.5 Binary Code Similarity Detection Ablation Study
Table $14$ show the detailed ablation study results of BCSD. Nova wins on the most benchmarks when $K=100$ or $500$, and ties with Nova-HA when $K=50$, or $200$.
A.6 Additional Analysis of Embedding
A.7 Additional Analysis of Attention

### v6

@@ 5.2 Binary Code Similarity Detection
Tables 9,  9,  9 and  9 show the Recall@1 of Nova and existing BCSD techniques with pool size $K$ of 50, 100, 200 and 500 on the six benchmarks. Underline indicates the best in each benchmark, while wave denotes the tied best (we only mark Nova-1.3B for clearer illustration).
Overall, Tables 9,  9,  9 and 9 show that on average, Nova-1.3B and Nova-6.7B achieve the highest Recall@1 (in bold) under all four settings of $K$. Nova-6.7B further outperforms Nova-1.3B and achieves the highest averaged Recall@1 under all four settings, ranking the ground-truth of 5%, 2%, 4%, and 3% more queries the most similar correspondingly compared to CodeArt. Nova-1.3B consistently outperforms existing techniques with higher Recall@1 when $K$ is 50, 100, and 200, meaning it correctly ranks ground-truth of 3%, 1%, and 2% more queries as the most similar. Under the setting of $K=500$, Nova-1.3B ties with CodeArt with the same highest Recall@1. When looking into each individual benchmark, Nova-1.3B always wins on the most benchmarks under different settings of pool size $K$. For instance, Nova-1.3B wins on four benchmarks while DiEmph only wins on SQLite when $K=50$. We also conduct an ablation study on BCSD in Appendix A.6.
@@ 5.3 Analytic Experiments: How are Nova’s embeddings better?
Figure 5: t-SNE analysis of embeddings calculated by Nova-CL-HA, Nova-HA, and Nova.
@@ 6 Conclusion
This work presents Nova, a generative foundation LLM for assembly code that introduces two key innovations—hierarchical attention and contrastive learning—to tackle the unique challenges of assembly code understanding. Nova builds upon a source code LLM and undergoes further pre-training on approximately 4.3 million collected assembly functions. We fine-tune and evaluate Nova on two downstream tasks: binary code decompilation and binary code similarity detection. Nova achieves up to 14.84 – 21.58% (absolute percentage point improvement) higher Pass@1 and Pass@10 over existing methods in binary code decompilation and up to a 6.17% higher Recall@1 in binary code similarity detection. Looking forward, we believe that the proposed hierarchical attention and contrastive learning methods hold promise for enhancing foundation models in both source code and natural language domains, which we leave as future work.
@@ ACKNOWLEDGMENTS
This research was supported in part by NSF 1901242 and 2006688, a CFI fund, and J.P. Morgan AI Faculty Research Awards. Any opinions, findings, and conclusions in this paper are those of the authors only and do not necessarily reflect the views of our sponsors.
@@ A.1 Data Collection
There may be alternative ways of collecting assembly code, e.g., using gcc -S to directly dump the assembly code without producing executables. However, a key difference is that the assembly generated by gcc -S does not undergo the linking step. In practice, binary decompilation and analysis are typically performed on executable or linked assembly code, as it includes linker modifications and reflects the final binary structure. Our data collection aligns better with the practical use of assembly. Nevertheless, Nova’s approaches should be generalizable to assembly code obtained with a different approach, and the experiments remain as future work.
@@
A.2 Experimental Setup Details
This section provides additional details of training. We pre-train Nova starting from DeepSeek-Coder, using the language modeling objective ($L_{lm}$) for one epoch on the C functions and assembly functions collected from AnghaBench and The-Stack corpora. This step uses a batch size of 128, with the input truncated by a 1,024 tokens limit. The model weights are updated using the AdamW optimizer. The learning rate is $5e^{-5}$, using 1000 steps of warm-up and a cosine decay to adjust the learning rate.
@@ A.2 Experimental Setup Details
For the evaluation of BCSD, we reuse the framework provided by CodeArt Su et al. (2024) to evaluate the binary code similarity detection results once Nova produces the embeddings for functions in the test dataset. For each one of the $K$ functions, the O0 assembly function is used as the query to calculate the cosine similarity between its embedding and the embeddings of all the $K$ O3 assembly functions. Note that Nova does not normalize the embeddings of assembly functions during training and uses $l_{2}$ distance to calculate the $L_{BCSD}$, which optimize Nova to embed the O0 and O3 assembly functions from the same source code have the smallest $l_{2}$ distances. When using CodeArt’s framework for evaluation which ranks the similarity using the cosine similarity, we normalize Nova’s embedding for each assembly function since $l_{2}$ distance after normalization keeps the same order as cosine similarity (smaller $l_{2}$ distance means higher cosine similarity).
Infrastructure The experiments are conducted on eight NVIDIA RTX A100 GPUs, each with 40GB memory. Our implementation is based on Huggingface’s implementation of DeepSeek-Coder  [^1] https://huggingface.co/deepseek-ai/deepseek-coder-1.3b-base, PyTorch  [^2] https://pytorch.org/get-started/locally/, and DeepSpeed  [^3] https://github.com/microsoft/DeepSpeed.
@@ A.3 Applying Hierarchical Attention on Half Attention Heads
Table 11: Ablation study of Nova-1.3B on HumanEval-Decompile.
Techniques Pass@1 Pass@10
O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
Nova-FCL-HA 22.38 16.20 16.37 13.25 17.05 36.13 29.48 30.02 23.76 29.85
Nova-OCL-HA 28.44 18.87 18.53 15.76 20.40 40.28 32.33 31.80 27.05 32.87
Nova-HA 30.58 19.88 20.58 16.40 21.86 44.75 33.13 33.31 29.82 35.25
Nova 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
A.4 Additional Ablation Study on Binary Code Decompilation
We provide additional ablation studies on studying the impact of each individual contrastive learning objective. We study two more models:
  • Nova-FCL-HA: Removing functional contrastive learning and hierarchical self-attention.
  • Nova-OCL-HA: Removing optimization contrastive learning and hierarchical self-attention.
Table 11 shows that each component of the contrastive learning brings certain improvements to the Pass@1 and Pass@10 on HumanEval-Decompile, and we find the impact of function contrastive learning (FCL) is larger than the impact of optimization contrastive learning (OCL), suggesting that aligning the model’s embeddings for assembly code with the same functionality is more useful.
Figure 7: Nova-1.3B correctly decompiles HumanEval-Decompile task 0.
Figure 8: Nova-1.3B correctly decompiles HumanEval-Decompile task 20.
A.5 Binary Code Decompilation Case Studies
Figure 7 shows an example from HumanEval-Decompile (task_id 0). Given the O1-optimized assembly code, GPT-4 fails to figure out the number of function arguments correctly, missing one important argument “float e”, and thus produces wrong functionality in the decompiled code. LLM4Decompile-1.3B makes similar mistakes and also misses the inner nested for loop. Nova-1.3B correctly decompiles the assembly into source code, where the ground truth is checking if any two elements in the given list *a (with size n) are close to each other than a given threshold e.
Figure 8 shows another more complex example, HumanEval-Decompile task_id 20. Nova-1.3B correctly decompiles the source code, successfully figuring that the function is trying to find the two elements that are closest to each other in the given array *a, with the result stored in minmax.
Table 12: Ablation study with $K=50$.
Benchmarks Nova-CL-HA Nova-HA Nova-1.3B
@@ A.5 Binary Code Decompilation Case Studies
Table 13: Ablation study with $K=100$.
Benchmarks Nova-CL-HA Nova-HA Nova-1.3B
@@
Table 14: Ablation study with $K=200$.
Benchmarks Nova-CL-HA Nova-HA Nova-1.3B
@@
Table 15: Ablation study with $K=500$.
Benchmarks Nova-CL-HA Nova-HA Nova-1.3B
@@
A.6 Binary Code Similarity Detection Ablation Study
Table $15$ show the detailed ablation study results of BCSD. Nova wins on the most benchmarks when $K=100$ or $500$, and ties with Nova-HA when $K=50$, or $200$.
@@ A.7 Additional Analysis of Embedding
A.8 Additional Analysis of Attention
@@
A.9 Limitations
One limitation is that Nova is X86-specific, as we only collect X86 assembly corpus for pre-training. This design choice is mainly affected by two facts: (1) X86 assembly is used and explored in a wide range of binary tasks (Wang et al., 2022; Su et al., 2024; Xu et al., 2023a; Chen et al., 2022c) compared to other assembly languages, and (2) computation limitations. However, the proposed techniques are independent of X86 assembly. Low information density and compiler optimization are the common challenges of most assembly languages such as X86, ARM, and MIPS. The proposed techniques can be applied to the future development of multi-lingual assembly LLMs. Another potential limitation is the scale of models. We develop Nova-1.3B and Nova-6.7B and demonstrate their advancement in assembly code decompilation and encoding. Developing larger Nova models is promising, but remains as future work due to limited computing resources.